Now suppose that Tommy is very adept at avoiding cars but will be injured if a truck passes during his attempt to cross the highway. The proportion of trucks to cars on Highway 6 is approximately 1 to 4. Assume that vehicles arrive in accordance with a Poisson process with rate λ = 4 per minute, and that a given vehicle that arrives has a 0.2 probability of being a truck and a 0.8 probability of being a car. What is the probability that Tommy will remain unhurt if it takes him c seconds to cross? Calculate specific probabilities for c = 3,5, 15.
Answer #1

Tommy will remain unhurt if no vehicles arrive in time c or if any vehicles arrive all are cars.

The number of customers arriving in c seconds can be modelld by Poisson distribution.

The Poisson PMF is

P\left ( X=x \right )=e^{-\lambda c}\frac{\left (\lambda c \right )^x}{x!};x=0,1,2,3,..... Here \lambda =4/60=1/15,c=3,5,15

The probability of no vehicles is

P\left ( X=0 \right )=e^{-\lambda c}\frac{\left (\lambda c \right )^0}{0!}=e^{-\lambda c}

The probability of X vehicles in c seconds and all are cars is

0.8^xP\left ( X=x \right )=0.8^xe^{-\lambda c}\frac{\left (\lambda c \right )^x}{x!};x=1,2,3,....

Thus the probability of not being hit is

P\left ( \textup{unhurt} \right )=P\left ( X=0 \right )+\sum_{x=1}^{\infty }0.8^xe^{-\lambda c}\frac{\left (\lambda c \right )^x}{x!}\\ P\left ( \textup{unhurt} \right )= P\left ( X=0 \right )+\sum_{x=1}^{\infty }e^{-\lambda c}\frac{\left (0.8\lambda c \right )^x}{x!}\\ P\left ( \textup{unhurt} \right )=P\left ( X=0 \right )+e^{-\lambda c}\sum_{x=1}^{\infty }\frac{\left (0.8\lambda c \right )^x}{x!}\\ P\left ( \textup{unhurt} \right )=e^{-\lambda c}+e^{-\lambda c}\left ( e^{0.8\lambda c}-1 \right )\\ P\left ( \textup{unhurt} \right ) =e^{-0.2\lambda c}

For c=3,5,15 , the probabilities are

{\color{Blue} P\left (\textup{unhurt} \right )=0.9607894, 0.9355070 ,0.8187308}